OpenStudy (anonymous):

Consider a random walk on the discrete line in which the particle’s initial position is 0, and the position increases by 1 in a step with the probability p and decreases by 1 in a step with the probability q = 1 − p, independently over different time steps, where 0 < p < 1. Given an integer x > 0, what is the probability that the particle ever visits the point of coordinate x? I have no idea where to even begin. There is a mess of equations in the class notes, and he's offered a hint, but I'm unsure how they go together.

ganeshie8 (ganeshie8):

what's the hint ?

OpenStudy (anonymous):

Hm. Yeah, that may be helpful, huh. "We used the notation $S_1(1)$ for the probability the particle ever visits the region to the right. Moving to the right means moving from y to y + 1. To get from 0 to x, the particle must visit 0, 1, 2, ... x - 1, x"

OpenStudy (anonymous):

After some crazy arithmetic I don't understand, he says \[S_1(1)=\frac{p}{q}\]

OpenStudy (anonymous):

One of the "I don't understand" is \[\sqrt{1-4pq} = |p - q|\] .. Wat?

OpenStudy (anonymous):

crap. \[S_1(z)=\frac{1-(1-4pqz^2)^{1/2}}{2qz}\]

OpenStudy (anonymous):

To at least clarify one result, we have that \(\begin{aligned}\sqrt{1-4pq} &= \sqrt{1-4p(1-p)}\\&= \sqrt{1-4p+4p^2}\\ &= \sqrt{(1-2p)^2}\\ &= \sqrt{(1-p -p)^2}\\&= \sqrt{(q-p)^2} \\ &= |p-q| \end{aligned}\)

OpenStudy (anonymous):

Great. That makes sense! So obvious now.

ganeshie8 (ganeshie8):

I think you need to also specify how many steps are allowed, because if you allow infinitely many steps, the probability for reaching any x ccoordinate is 1

OpenStudy (anonymous):

Yeah. I wonder if that's what he's looking for.

ganeshie8 (ganeshie8):

Maybe you could find the probability for visiting the value "x" at "nth" step, then add the probabilities for \(n \in \mathbb{N}\)

OpenStudy (anonymous):

I believe probability would be more like \(\left(\dfrac{p}{q}\right)^x\) if you hit each one in order of 1,2,.., x-1, x without moving to the left once...but that would just be too easy. XD

OpenStudy (anonymous):

I'm not sure. This class drives me insane.

ganeshie8 (ganeshie8):

To reach the value "x" at "nth" step, the particle must take (n+x)/2 moves to right and (n-x)/2 moves to left that should allow you to cookup an expression for probability for particle to be at position x at nth step

ganeshie8 (ganeshie8):

since the events, moving left/right, are indpendent, we can simply add up all the different probabilities

OpenStudy (anonymous):

What the ... The hint says the particle would move in 1, 2, ..., x - 1, x in that order. But that doesn't make sense. if x = 4, it could go 1, 2, 3, 2, 3, 4. To your point @ganeshie8

ganeshie8 (ganeshie8):

right, there are infinitely many ways to reach the point x if you are not fixing the "number of steps taken"

OpenStudy (anonymous):

Alright. I think I have what I need. Thank you.

ganeshie8 (ganeshie8):

go through this if u can http://physics.gu.se/~frtbm/joomla/media/mydocs/LennartSjogren/kap2.pdf it derives below expression for the probability for the particle to be at coordinate m after N steps 

OpenStudy (anonymous):

Even better. Makes sense! Appreciate the help!